5 A store records number of costumers per hour and has the statistic below:

costumers/h 0 1 2 3 >3

probability 0.05 0.15 0.22 0.22 0.36

a) What is the probability the max 2 costumers come during an hour?

b) What is the probability the at least 3 costumers come during an hour?

c) What is the probability the at least some costumer come during an hour?
Answer #1

Ans:

a)P(x<=2)=P(x=0)+P(x=1)+P(x=2)

=0.05+0.15+0.22

=0.42

b)P(x>=3)=0.22+0.36=0.58

c)P(atleast some)=P(x>=1)=1-P(x=0)=1-0.05=0.95